# The Energy Cost of Artificial Intelligence of Things Lifecycle

### v2

@@ IV Energy cost of Data Collection
In this section, we derive the energy cost of data collection in an AI-powered wireless system. To calculate such cost, we first determine the number of bits that need to be transmitted from an IoT device to the server via wireless technology (uplink). Then, we calculate the energy consumption of both transmitting and receiving the data based on the given transmitting/ receiving power and corresponding transmission/ receiving rate. Finally, we derive the total energy consumption for collecting data from the IoT device. Note that the energy consumption of wired communication is considered negligible in comparison to the wireless part [31]. Therefore, this work focuses solely on the communication from the IoT device to the access point (AP) via wireless technologies, however the open source calculator can be extended to also include other communication segments and technologies.
@@ IV-A OSI Overhead
In OSI architecture, each layer introduces its own overhead, which includes various fields in the headers in the Data Plane (DP) and additional control and set-up messaging from the Control Plane (CP). Additionally, in some of the layers, such as link (e.g., MAC) and transport (e.g., TCP) layers, errors or packet loss are internally treated, potentially leading to retransmissions. While each protocol and every layer enables a wide selection of settings, we devise the following generalized theoretical formulation for computing the size of the transmitted data at the physical layer ($B_{\mathrm{\mathrm{T}}}$):
@@
$B_{\mathrm{T}}$ $\leavevmode\nobreak\ [b]=\lceil B_{\mathrm{T,L_{\mathrm{OSI}}}}\prod_{l=1}^{L_{\mathrm{OSI}}}\underbrace{\left(RR_{\mathrm{DP},l}\cdot(1+OH_{\mathrm{DP},l})\right.}_{\text{contribution from DP}}$
@@
$+\underbrace{\left.RR_{\mathrm{CP},l}\cdot\gamma_{l}\cdot OH_{\mathrm{CP},l}\right)}_{\text{contribution from CP}}\rceil,\mathrm{where}\leavevmode\nobreak\ L_{\mathrm{OSI}}=7,$ (2)
@@
where $B_{\mathrm{T,L_{OSI}}}$ represents the application layer data (i.e. IoT device measurements) sent from the IoT device, measured in bits. It is calculated as the product of the bit precision ($f$), the number of samples ($N_{\mathrm{S}}$) and the sample size ($I_{\mathrm{S}}$) in floating-point format, i.e., $B_{\mathrm{T,L_{\mathrm{OSI}}}}=fN_{\mathrm{S}}I_{\mathrm{S}}$. Typically, $f$ equals $32$ bits for single precision (float) or 64 bits for double precision format. To be able to send/receive data, in each of the seven OSI layers, represented by the product in Eq. 2, a number of data and control plane operations take place as expressed in the respective terms. The application data, $B_{\mathrm{T,L_{OSI}}}$, is processed and is added some overhead $OH_{\mathrm{DP},l}$, in some layers it is encrypted (i.e. application, presentation and/or transport layers) and in some layers it may also require retransmissions $RR_{\mathrm{DP},l}$. Moreover, some layers, such as the network layer, incur additional overhead required from signaling to maintain routing tables. This is reflected in the second term of Eq. 2 through the respective CP overheads $OH_{\mathrm{CP},l}$, retransmissions $RR_{\mathrm{CP},l}$ and the $\gamma_{l}$ scaling factor that adjusts for the relative contribution of the CP overhead compared to the DP counterpart at the $l$-th layer.
@@
The eCAL calculator implements the computation from Eq. 2 and enables the configuration of the application data size, retransmission rates, overheads, and scaling factor. Furthermore, it enables plugging custom protocol implementations for each of the layers, based also on the existing literature [32, 33, 34], so that the community can also instantiate eCAL for a specific well-defined set of protocols in addition to the generalized theoretical approach adopted in this paper.
@@
To illustrate the behavior of Eq. 2, we first consider a scenario where a single IoT device transmits $N_{\mathrm{S}}$ of double-precision samples to the application accompanied by DP and CP overheads while altering the retransmission rate. In particular, we consider that each layer in the OSI model introduces a fixed overhead percentage for the DP and CP, with $10$% and $5$%, respectively. Furthermore, the relative contribution of CP overhead is scaled by a factor of $0.5$, i.e., $\gamma_{l}=0.5$. The retransmission rates for both DP and CP are assumed to be equal and vary from $1$ to $2$ in increments of $0.04$. This is shown in Fig. 2. The results show the retransmission rate has a significant impact on total transmitting bits. For example, when $N_{\mathrm{S}}=256$ samples, there is a growth of $2.3$, $17$, and $128$ times with retransmission rates equal to $1.125$, $1.5$, and $2$, respectively, compared to scenarios with no retransmission.
@@ IV-B Energy Estimation
To estimate the energy consumption of the data collection component ($E_{\mathrm{DC}}$), which is the sum of the energy of transmitting ($E_{\mathrm{T}}$) and receiving the data ($E_{\mathrm{R}}$) introduced in the physical layer (PHY), as well as the energy consumption of computation in the DP and CP across the various layers of OSI ($E_{\mathrm{C}}$), we propose the following analytical expression:
$E_{\mathrm{DC}}$ $=\underbrace{\frac{P_{\mathrm{T}}[W]}{R_{\mathrm{T}}[b/s]}\cdot B_{\mathrm{T}}[b]}_{\text{PHY layer transmitting energy ($E_{\mathrm{T}}$)}}+\underbrace{\frac{P_{\mathrm{R}}[W]}{R_{\mathrm{R}}[b/s]}\cdot B_{\mathrm{T}}[b]}_{\text{PHY layer receiving energy ($E_{\mathrm{R}}$)}}$ (3)
$+\sum^{L_{\mathrm{OSI}}}_{l=2}(\underbrace{B_{\mathrm{T}_{l}}\leavevmode\nobreak\ [b]\cdot N_{\mathrm{IoT,cycle},l}\leavevmode\nobreak\ [c/b]\cdot P_{\mathrm{IoT,cycle}}\leavevmode\nobreak\ [J/c]}_{\text{from APP to MAC layer computational energy at the IoT device}}$
$+\underbrace{B_{\mathrm{T}_{l}}\leavevmode\nobreak\ [b]\cdot N_{\mathrm{gateway,cycle},l}\leavevmode\nobreak\ [c/b]\cdot P_{\mathrm{gateway,cycle}}\leavevmode\nobreak\ [J/c]}_{\text{from MAC to APP layer computational energy at the gateway}})$
@@
where $P_{\mathrm{T}}$ and $R_{\mathrm{T}}$ represent the transmitting power and data rate, respectively, while $P_{\mathrm{R}}$ and $R_{\mathrm{R}}$ are the receiving power and data rate. Furthermore, $B_{\mathrm{T},l}$ denotes the total number of the bits that need to be processed at the $l$-th layer in the OSI model. Additionally, we define the computational parameters at the IoT device as $N_{\mathrm{IoT,cycle},l}$ representing the number of processing cycles per bit at the $l$-th layer, and $P_{\mathrm{IoT,cycle}}$, denoting the power consumption per cycle. Similarly, at the gateway, these parameters are defined as $N_{\mathrm{gateway,cycle},l}$ and $P_{\mathrm{gateway,cycle}}$.
@@
Computing exact values for $E_{\mathrm{DC}}$ for a broad range of $B_{\mathrm{T}_{l}}$, $N_{\mathrm{\_,cycle},l}$ and $P_{\mathrm{\_,cycle}}$ is challenging due to the overall complexity of computing systems. For instance, for processing the same amount of data, $N_{\mathrm{\_,cycle,}l}$ varies across microprocessors while also not having a linear dependency with the data size [35]. Therefore, the values for $B_{\mathrm{T}_{l}}$ and $N_{\mathrm{\_,cycle}}$ cannot be chosen arbitrarily. Next, $P_{\mathrm{\_,cycle}}$ also depends on the microprocessor. However, in Fig. 3, we illustrate the energy consumption of $E_{\mathrm{DC}}$ with a selected set of values to provide an intuition on the contribution from each term. More specifically, we analyze a scenario where $N_{\mathrm{S}}=256$, and we consider two sets of parameters for computing transmitting and receiving energy, i.e. $E_{\mathrm{T}}$ and $E_{\mathrm{R}}$, alongside two different settings for calculating the computational energy consumption ($E_{\mathrm{C}}$). The results show that while transmission energy ($E_{\mathrm{T}}$) remains the dominant contributor in most cases, the computational energy consumption ($E_{\mathrm{C}}$) is not negligible and, in some configurations, constitutes a significant portion of the total energy consumption of data collection. This is particularly the case when the hardware is less optimized for the tasks [35], leading to higher cycles per bit ($N_{\mathrm{\_,cycle},l}$) and power per cycle ($P_{\mathrm{\_,cycle}}$). For example, when $N_{\mathrm{\_,cycle}}=100$ and $P_{\mathrm{\_,cycle}}=2\times 10^{-10}$, $E_{\mathrm{C}}$ is $7.55\times 10^{-3}\leavevmode\nobreak\ [J]$, which represents approximately $10\%$ and $95\%$ of the total energy cost in two different PHY layer configurations. On the other hand, when $N_{\mathrm{\_,cycle},l}=50$ and $P_{\mathrm{\_,cycle}}=10^{-10}$, $E_{\mathrm{C}}$ is $6.61\times 10^{-3}\leavevmode\nobreak\ [J]$, contributing to approximately $5\%$ and $90\%$ of the total energy consumption in data collection component.
@@ V Energy Cost of Data Preprocessing
At the server side, data is first stored on a hard drive, subsequently preprocessed, and then used for the ML model development or inference, but we are only focusing on the latter in Section VII. While storage is generally energy-efficient and in terms of energy costs relatively negligible, it can become significant when handling large volumes of incoming data or when utilizing virtualized hosted storage with encryption in an edge-cloud environment [36]. In the following, we consider the dataset with $N_{\mathrm{S}}$ samples collected via the wireless network, as depicted in Fig. 1 and discussed in Section IV.
In this study, we consider data preprocessing consisting of data cleaning, and data transformation steps, while leaving other more complex approaches, such as feature engineering and feature for future work. In the data cleaning step, the task of preprocessing is to identify and remove invalid samples from a list or array which incurs zero floating-point operations (FLOPs) since no floating points additions and multiplications are performed, only comparisons and memory operations. In the data transformation step, the preprocessing needs to determine the complexity of data transformation ($M_{\mathrm{DS}}$). We provide three examples, i.e., min-max scaling, normalization, and Gramian Angular Difference Field (GADF) process which represent typical approaches for data transformation.
@@
Min-max scaling is a data preprocessing technique that transforms numerical data to fit within the [0,1] interval by scaling each value relative to the minimum and maximum values in the dataset. For each feature, the transformation subtracts the minimum value and divides by the range, ensuring that the lowest value becomes 0 and the highest value becomes 1, with all other values scaled proportionally between them. While finding the minimum and maximum values requires computational resources for comparisons, these operations do not contribute to the FLOPs count. The scaling itself requires calculating the range (one FLOP), followed by subtracting the minimum and dividing by the range for each sample ($2\cdot N_{\mathrm{S}}\cdot I_{\mathrm{S}}$ FLOPs). Therefore, the total number of FLOPs ($M^{\mathrm{minmax}}_{\mathrm{DS}}$) for min-max scaling can be expressed as:
$M^{\mathrm{minmax}}_{\mathrm{DS}}=2\cdot N_{\mathrm{S}}\cdot I_{\mathrm{S}}+1,$ (4)
@@
Standard score, also known as z-score normalization, represents a more computationally intensive preprocessing technique compared to min-max scaling. The process involves three distinct computational steps. First, calculating the population mean requires $N_{\mathrm{S}}\cdot I_{\mathrm{S}}$ FLOPs, as we sum all values and divide by the total number of samples. Second, computing the population standard deviation demands $13\cdot N_{\mathrm{S}}\cdot I_{\mathrm{S}}+$ 1 FLOPs, which encompasses calculating squared differences from the mean, their sum, and the final square root operation. Finally, the standardization transformation itself necessitates $2\cdot N_{\mathrm{S}}\cdot I_{\mathrm{S}}$ FLOPs, as each data point must be centered by subtracting the mean and scaled by dividing with the standard deviation. The total computational cost for normalization, expressed in FLOPs ($M^{\mathrm{norm}}_{\mathrm{DS}}$), is therefore:
$M^{\mathrm{norm}}_{\mathrm{DS}}=6\cdot N_{\mathrm{S}}\cdot I_{\mathrm{S}}+1$ (5)
@@
The computational complexity of GADF can be broken down into its constituent operations. The initial min-max scaling requires $12\cdot N_{\mathrm{S}}\cdot I_{\mathrm{S}}+$ 1 FLOPs as previously discussed, however in these case we have to multiply the number of samples with the length of the input time series ($I_{\mathrm{S}}$) as we are dealing with a series of values in a single sample increasing the FLOPs. The polar coordinate transformation and the construction of the GADF matrix itself, requires $(5\cdot I_{\mathrm{S}}+I_{\mathrm{S}}^{2})\cdot N_{\mathrm{S}}$ FLOPs . Therefore, the total number of FLOPs ($M^{\mathrm{GADF}}_{\mathrm{DS}}$) for the GADF transformation can be expressed as:
$M^{\mathrm{GADF}}_{\mathrm{DS}}=\left(2\cdot N_{\mathrm{S}}\cdot I_{\mathrm{S}}+1\right)+(5\cdot I_{\mathrm{S}}+I_{\mathrm{S}}^{2})\cdot N_{\mathrm{S}}$ (6)
@@
$M_{\text{pre}}=\begin{cases}2\cdot N_{\text{S}}\cdot I_{\mathrm{S}}+1,\text{min-max scaling},\\ 6\cdot N_{\text{S}}\cdot I_{\mathrm{S}}+1,\text{normalization},\\ 2\cdot N_{\mathrm{S}}\cdot I_{\mathrm{S}}+1+(5\cdot I_{\mathrm{S}}+I_{\mathrm{S}}^{2})\cdot N_{\mathrm{S}},\text{GADF}.\end{cases}$ (7)
@@
$E_{\mathrm{pre}}[J]=P_{\mathrm{pre}}[W]\cdot T_{\mathrm{pre}}[s],$ (8)
@@
Figure 5: Energy consumption ($E_{\mathrm{pre}}$) of fixed data point ($N_{\mathrm{S}}\cdot I_{\mathrm{S}}$) across different preprocessing techniques and sample size ($I_{\mathrm{S}}$).

### v3

@@ IV Energy cost of Data Collection
In this section, we derive the energy cost of data collection in an AI-powered wireless system. To calculate this cost, we first determine the number of bits that need to be transmitted from an device to the server via wireless technology (uplink). Then, we calculate the energy consumption of both transmitting and receiving the data, based on the given transmission and reception power and the corresponding rate. Finally, we derive the total energy consumption for collecting data from the device. It is important to note that the energy consumption of wired communication is considered negligible in comparison to the wireless part [31]. Therefore, this work focuses solely on the communication from the wireless device to the access point (AP) via wireless technologies. The assumption is that all data is collected via a single link; however, a summation can be added for multiple links. The modular and extensible open source calculator can be extended to include specific communication segments, topologies and technologies employing various resource allocation and optimization techniques.
@@ IV-A OSI Overhead
In OSI architecture, each layer introduces its own overhead, including various header fields in the Data Plane (DP) and additional control and set-up messaging from the Control Plane (CP). Additionally, in some of the layers, such as link (e.g., MAC) and transport (e.g., TCP) layers, errors or packet loss are internally treated, potentially leading to retransmissions. While each protocol and every layer enables a wide selection of settings, we devise the following generalized theoretical formulation for computing the size of the transmitted data at the physical layer ($B_{\mathrm{\mathrm{T}}}$):
@@
$B_{\mathrm{T}}$ $\leavevmode\nobreak\ [b]=\lceil B_{\mathrm{T,L_{\mathrm{OSI}}}}\prod_{l=1}^{L_{\mathrm{OSI}}}\underbrace{\left(RR_{\mathrm{DP},l}(1+OH_{\mathrm{DP},l})\right.}_{\text{contribution from DP}}$
@@
$+\underbrace{\left.RR_{\mathrm{CP},l}\gamma_{l}OH_{\mathrm{CP},l}\right)}_{\text{contribution from CP}}\rceil,\leavevmode\nobreak\ \mathrm{where}\leavevmode\nobreak\ L_{\mathrm{OSI}}=7,$ (2)
@@
where $B_{\mathrm{T,L_{OSI}}}$ represents the application layer data (i.e. telemetry) sent from the device, measured in bits. It is calculated as the product of the bit precision ($f$), the number of samples ($N_{\mathrm{S}}$) and the sample size ($I_{\mathrm{S}}$) in floating-point format, i.e., $B_{\mathrm{T,L_{\mathrm{OSI}}}}=fN_{\mathrm{S}}I_{\mathrm{S}}$. Typically, $f$ equals $32$ bits for single precision (float) or 64 bits for double precision format. To be able to send/receive data, in each of the seven OSI layers, represented by the product in Eq. 2, a number of data and control plane operations take place as expressed in the respective terms. The application data, $B_{\mathrm{T,L_{OSI}}}$, is processed and is added some overhead $OH_{\mathrm{DP},l}$, in some layers it is encrypted (i.e. application, presentation and/or transport layers) and in some layers it may also require retransmissions $RR_{\mathrm{DP},l}$. Moreover, some layers, such as the network layer, incur additional overhead required from signaling to maintain routing tables. This is reflected in the second term of Eq. 2 through the respective CP overheads $OH_{\mathrm{CP},l}$, retransmissions $RR_{\mathrm{CP},l}$ and the $\gamma_{l}$ scaling factor that adjusts for the relative contribution of the CP overhead compared to the DP counterpart at the $l$-th layer. The scaling factor can be also used to compensate for other specific overheads unforeseen by the considered configuration in this paper.
@@
The eCAL calculator implements the computation from Eq. 2 and enables the configuration of the application data size, retransmission rates, overheads, and scaling factor. Furthermore, it enables plugging custom protocol implementations for each of the layers, based also on the existing literature [32, 33, 34], so that the community can also instantiate eCAL for a specific well-defined set of protocols in addition to the approach adopted in this paper.
@@
To illustrate the behavior of Eq. 2, we first consider a scenario where a single device that transmits $N_{\mathrm{S}}$ of double-precision samples to the application accompanied by DP and CP overheads while altering the retransmission rate. In particular, we consider that each layer in the OSI model introduces a fixed overhead percentage for the DP and CP, with $10$% and $5$%, respectively. Furthermore, the relative contribution of CP overhead is scaled by a factor of $0.5$, i.e., $\gamma_{l}=0.5$. The retransmission rates for both DP and CP are assumed to be equal and vary from $1$ to $2$ in increments of $0.04$. This is shown in Fig. 2. The results show the retransmission rate has a significant impact on total transmitting bits. For example, when $N_{\mathrm{S}}=256$ samples, there is a growth of $2.3$, $17$, and $128$ times with retransmission rates equal to $1.125$, $1.5$, and $2$, respectively, compared to scenarios with no retransmission.
@@ IV-B Energy Estimation
To estimate the energy consumption of the data collection component ($E_{\mathrm{DC}}$), which is the sum of the energy of transmitting ($E_{\mathrm{T}}$) and receiving the data ($E_{\mathrm{R}}$) introduced in the physical layer (PHY), as well as the energy consumption of computation in the DP and CP across the various layers of OSI ($E_{\mathrm{C}}$). We propose the following analytical expression:
$E_{\mathrm{DC}}$ $=\underbrace{\frac{P_{\mathrm{T}}[W]}{R_{\mathrm{T}}[b/s]}B_{\mathrm{T}}[b]}_{\text{PHY layer transmitting energy ($E_{\mathrm{T}}$)}}+\underbrace{\frac{P_{\mathrm{R}}[W]}{R_{\mathrm{R}}[b/s]}B_{\mathrm{T}}[b]}_{\text{PHY layer receiving energy ($E_{\mathrm{R}}$)}}$ (3)
$+\sum^{L_{\mathrm{OSI}}}_{l=2}(\underbrace{B_{\mathrm{T}_{l}}\leavevmode\nobreak\ [b]N_{\mathrm{dev,cycle},l}\leavevmode\nobreak\ [c/b]P_{\mathrm{dev,cycle}}\leavevmode\nobreak\ [J/c]}_{\text{from APP to MAC layer computational energy at the device}}$
$+\underbrace{B_{\mathrm{T}_{l}}\leavevmode\nobreak\ [b]N_{\mathrm{gateway,cycle},l}\leavevmode\nobreak\ [c/b]P_{\mathrm{gateway,cycle}}\leavevmode\nobreak\ [J/c]}_{\text{from MAC to APP layer computational energy at the gateway}})$
@@
where $P_{\mathrm{T}}$ and $R_{\mathrm{T}}$ represent the transmitting power and the data rate, respectively, while $P_{\mathrm{R}}$ and $R_{\mathrm{R}}$ are the receiving power and data rate. Furthermore, $B_{\mathrm{T},l}$ denotes the total number of the bits that need to be processed at the $l$-th layer in the OSI model. Additionally, we define the computational parameters at the device as $N_{\mathrm{dev,cycle},l}$ representing the number of processing cycles per bit at the $l$-th layer, and $P_{\mathrm{dev,cycle}}$, denoting the power consumption per cycle. Similarly, at the gateway, these parameters are defined as $N_{\mathrm{gateway,cycle},l}$ and $P_{\mathrm{gateway,cycle}}$.
@@
Computing exact values for $E_{\mathrm{DC}}$ for a broad range of $B_{\mathrm{T}_{l}}$, $N_{\mathrm{\_,cycle},l}$, and $P_{\mathrm{\_,cycle}}$ is challenging due to the overall complexity of computing systems. For instance, $N_{\mathrm{\_,cycle,}l}$ varies across microprocessors when processing the same amount of data, and it does not have a linear dependency with data size [35]. Therefore, the values for $B_{\mathrm{T}_{l}}$ and $N_{\mathrm{\_,cycle}}$ cannot be chosen arbitrarily. Next, $P_{\mathrm{\_,cycle}}$ also depends on the microprocessor. However, in Fig. 3, we illustrate the energy consumption of $E_{\mathrm{DC}}$ with a selected set of values to provide an intuition on the contribution from each term. More specifically, we analyze a scenario where $N_{\mathrm{S}}=256$, and we consider two sets of parameters for computing transmitting and receiving energy, i.e. $E_{\mathrm{T}}$ and $E_{\mathrm{R}}$, alongside two different settings for calculating the computational energy consumption ($E_{\mathrm{C}}$). The results show that while transmission energy ($E_{\mathrm{T}}$) remains the dominant contributor in most cases, the computational energy consumption ($E_{\mathrm{C}}$) is not negligible and, in some configurations, constitutes a significant portion of the total energy consumption of data collection. This is particularly the case when the hardware is less optimized for the tasks [35], leading to higher cycles per bit ($N_{\mathrm{\_,cycle},l}$) and power per cycle ($P_{\mathrm{\_,cycle}}$). For example, when $N_{\mathrm{\_,cycle}}=100$ and $P_{\mathrm{\_,cycle}}=2\times 10^{-10}$, $E_{\mathrm{C}}$ is $7.55\times 10^{-3}\leavevmode\nobreak\ [J]$, which represents approximately $10\%$ and $95\%$ of the total energy cost in two different PHY configurations. On the other hand, when $N_{\mathrm{\_,cycle},l}=50$ and $P_{\mathrm{\_,cycle}}=10^{-10}$, $E_{\mathrm{C}}$ is $6.61\times 10^{-3}\leavevmode\nobreak\ [J]$, contributing to approximately $5\%$ and $90\%$ of the total energy consumption in data collection component.
@@ V Energy Cost of Data Preprocessing
At the server side, data is first stored on a hard drive, subsequently preprocessed, and finally used for the ML model development or inference. This section focuses on preprocessing, while inference is addressed in Section VII. Storage is generally energy-efficient, with relatively negligible energy costs. However, it can become significant when handling large volumes of incoming data or when utilizing virtualized hosted storage with encryption in an edge-cloud environment [36]. In the following case study, we consider the dataset with $N_{\mathrm{S}}$ samples collected via the wireless network, as depicted in Fig. 1 and discussed in Section IV. We also consider data preprocessing consisting of data cleaning, and data transformation steps, while leaving other more complex approaches, such as feature engineering, for further extensions of the study and simulator.
In the data cleaning step, the task of preprocessing is to identify and remove invalid samples from a list or array, which incurs zero floating-point operations (FLOPs [^3] FLOPs are often use to enable theoretical estimation of computational complexity. Alternative approaches would involve measurements through interfaces such as PAPI or tools such as Scaphandre when measuring directly energy., as no floating-point additions or multiplications are performed, only comparisons and memory operations. During the data transformation step, preprocessing involves determining the complexity of data transformation ($M_{\mathrm{DS}}$). We provide three examples, i.e., min-max scaling, normalization, and Gramian Angular Difference Field (GADF) process which represent typical approaches for data transformation.
@@
Min-max scaling is a data preprocessing technique that transforms numerical data to fit within the [0,1] interval by scaling each value relative to the minimum and maximum values in the dataset. For each feature, the transformation subtracts the minimum value and divides by the range, ensuring that the lowest value becomes 0 and the highest value becomes 1, with all other values scaled proportionally between them. While finding the minimum and maximum values requires computational resources for comparisons, these operations do not contribute to the FLOPs count. The scaling itself requires calculating the range (one FLOP), followed by subtracting the minimum and dividing by the range for each sample ($2N_{\mathrm{S}}I_{\mathrm{S}}$ FLOPs). Therefore, the total number of FLOPs ($M^{\mathrm{minmax}}_{\mathrm{DS}}$) for min-max scaling can be expressed as:
$M^{\mathrm{minmax}}_{\mathrm{DS}}=2N_{\mathrm{S}}I_{\mathrm{S}}+1.$ (4)
@@
Standard score, also known as z-score normalization, represents a more computationally intensive preprocessing technique compared to min-max scaling. The process involves three distinct computational steps. First, calculating the population mean requires $N_{\mathrm{S}}I_{\mathrm{S}}$ FLOPs, as we sum all values and divide by the total number of samples. Second, computing the population standard deviation demands $13N_{\mathrm{S}}I_{\mathrm{S}}+$ 1 FLOPs, which encompasses calculating squared differences from the mean, their sum, and the final square root operation. Finally, the standardization transformation itself necessitates $2N_{\mathrm{S}}I_{\mathrm{S}}$ FLOPs, as each data point must be centered by subtracting the mean and scaled by dividing by the standard deviation. The total computational cost for normalization, expressed in FLOPs ($M^{\mathrm{norm}}_{\mathrm{DS}}$), is therefore:
$M^{\mathrm{norm}}_{\mathrm{DS}}=6N_{\mathrm{S}}I_{\mathrm{S}}+1.$ (5)
@@
The computational complexity of GADF can be broken down into its constituent operations. The initial min-max scaling requires $12N_{\mathrm{S}}I_{\mathrm{S}}+$ 1 FLOPs as previously discussed, however in these case we have to multiply the number of samples with the length of the input time series ($I_{\mathrm{S}}$) as we are dealing with a series of values in a single sample increasing the FLOPs. The polar coordinate transformation and the construction of the GADF matrix itself, requires $(5I_{\mathrm{S}}+I_{\mathrm{S}}^{2})N_{\mathrm{S}}$ FLOPs . Therefore, the total number of FLOPs ($M^{\mathrm{GADF}}_{\mathrm{DS}}$) for the GADF transformation can be expressed as:
$M^{\mathrm{GADF}}_{\mathrm{DS}}=\left(2N_{\mathrm{S}}I_{\mathrm{S}}+1\right)+(5I_{\mathrm{S}}+I_{\mathrm{S}}^{2})N_{\mathrm{S}}.$ (6)
@@
$M_{\text{pre}}=\begin{cases}2N_{\text{S}}I_{\mathrm{S}}+1,\text{min-max scaling},\\ 6N_{\text{S}}I_{\mathrm{S}}+1,\text{normalization},\\ 2N_{\mathrm{S}}I_{\mathrm{S}}+1+(5I_{\mathrm{S}}+I_{\mathrm{S}}^{2})N_{\mathrm{S}},\text{GADF}.\end{cases}$ (7)
@@
$E_{\mathrm{pre}}[J]=P_{\mathrm{pre}}[W]T_{\mathrm{pre}}[s],$ (8)
@@
Figure 5: Energy consumption ($E_{\mathrm{pre}}$) of fixed data point ($N_{\mathrm{S}}I_{\mathrm{S}}$) across different preprocessing techniques and sample size ($I_{\mathrm{S}}$).